OpenStudy (17jeisaac):

The area of the sector of a circle with a radius of 8 centimeters is 125.6 square centimeters. The estimated value of π is 3.14. The measure of the angle of the sector is

OpenStudy (shamim):

U know full angle of a circle is=360degree

OpenStudy (shamim):

Let the angle makes by ur given sector is= theta

OpenStudy (shamim):

Do u know area of s circle is=pi*r^2

OpenStudy (17jeisaac):

@shamim Yes

OpenStudy (shamim):

Given in ur question radius r=4cm Right?

OpenStudy (shamim):

Area of ur circle is=pi*r^2

OpenStudy (shamim):

=3.14*4^2=?

OpenStudy (17jeisaac):

@shamim Yes

OpenStudy (shamim):

Response plz!!!

OpenStudy (17jeisaac):

@shamim 25.12

OpenStudy (17jeisaac):

@shamim wait 50.24

OpenStudy (shamim):

Required formula is 360/theta=50.24/12.56 Theta=?